# Propagation speeds of relativistic conformal fluids from a generalized relaxation time approximation

### v1

@@
Abstract
We compute the propagation speeds for a conformal real relativistic fluid. We begin from a kinetic equation in the relaxation time approximation, where the relaxation time is an arbitrary function of the particle energy in the Landau frame. We propose a parameterization of the one particle distribution function designed to contain a second order Chapman-Enskog solution as a particular case. We derive the hydrodynamic equations applying the moments method to this parameterized one particle distribution function, and solve for the propagation speeds of linearized scalar, vector and tensor perturbations. For relaxation times of the form $\tau=\tau_{0}(-\beta_{\mu}p^{\mu})^{-a}$, with $-\infty<a<2$, where $\beta_{\mu}=u_{\mu}/T$ is the temperature vector in the Landau frame, we show that the Anderson-Witting prescription $a=1$ yields the fastest speeds.
I Introduction
The success of hydrodynamics in the description of the early stages of relativistic heavy ion collisions [1, 2] and the promise of relevant cosmological applications [3, 4, 5] have turned relativistic real fluids into an active area of research [6, 7]. However, no single set of hydrodynamic equations has been universally accepted as the proper description of these systems. In particular, attempts to derive hydrodynamics purely from macroscopic considerations [8, 9, 10, 11] have proven not compelling enough. For this reasons, the usual path to hydrodynamics is to derive it from a more microscopic theory, usually kinetic [12, 13, 14, 15] or field theory [16, 17]. In this paper, we shall adopt kinetic theory as the fundamental description, and derive hydrodynamic equations from it.
Given the complexity of nonlinear collision operators such as Boltzmann’s, oftentimes it is preferred to work under the Relaxation Time Approximation (RTA), which assumes a linear kinetic equation parametrized by a relaxation time. An early implementation of the RTA is Marle’s [18, 19], which assumes a single, momentum independent relaxation time. This was later improved on by Anderson and Witting [20, 21], which assumed the relaxation time was inversely proportional to the energy of the particle in the Landau frame of the fluid. However successful, both Marle and Anderson-Witting RTA’s severely distort the structure of the linearized Boltzmann equation, and their validity is doubtful for soft collision terms, where the continuous spectrum of the linearized collision operator reaches up to the zero eigenvalue associated with the hydrodynamic modes.
@@ I Introduction
This has led to more general implementations of the RTA [22, 23] where different modes were allowed to relax at different rates. However, simply allowing the relaxation times to be a function of energy other than constant or linear is not satisfactory as it may violate energy-momentum conservation. Rather the RTA must be implemented preserving the Hilbert space structure of the space of linearized one-particle distribution function (1-pdf’s). A kinetic equation allowing for an momentum dependent relaxation time consistent with energy-momentum conservation was introduced in [24, 25]. In this paper we shall elaborate on this proposal. In particular, we shall show how to produce a RTA matching any prescribed spectrum for the linearized collision operator, either soft or hard. Not been able to do this is one of the main drawbacks of the usual formulations of the RTA. See also [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]
The ultimate goal of this paper is to compute the propagation speeds for a conformal real relativistic fluid [38, 39]. To derive the hydrodynamic equations we apply the moments method to a kinetic equation under the RTA. We generalize the usual RTA by allowing an arbitrary dependence of relaxation time on the particle energy in the Landau frame of the fluid [24, 25]. We then find the propagation speeds of linearized scalar, vector and tensor perturbations. For relaxation times of the form $\tau=\tau_{0}(-\beta_{\mu}p^{\mu})^{-a}$, with $-\infty<a<2$, where $\beta_{\mu}=u_{\mu}/T$ is the temperature vector in the Landau frame, we show that the Anderson-Witting prescription $a=1$ yields the fastest speeds.
Given the kinetic equation, the following step is to derive hydrodynamics from it. We shall adopt the moments method, where one assumes a parameterized form for the 1pdf, and then derives equations for the parameters by taking moments of the kinetic equation. A rather general parameterization has the form [40]
$f=f_{0}e^{\chi}$ (1)
where $f_{0}$ is an equilibrium, Maxwell-Jüttner distribution (defined below, see eq. 3) and
$\chi=\sum_{\alpha}C_{\alpha}\left(x\right)X^{\alpha}\left(x,p\right)$ (2)
where the $X^{\alpha}$ are given functions of momentum, and the $C_{\alpha}$ are the parameters.
Any hydrodynamic approach based on kinetic theory faces the challenge of finding a suitable parameterization for the 1-pdf (in this paper, choosing the $X^{\alpha}$ functions) . The parameterization should be general enough to allow for an accurate description of physically meaningful processes but not so general as to make the ensuing theory unwieldly. Many proposals have been advanced in the literature [41, 42] . In this paper we shall adopt the point of view that the parameterization 1 and 2 must be such as to include the Chapman-Enskog (Ch-E) solution to the kinetic equation as a particular case [43]. This yields parameterizations with an increasing number of parameters depending on to which order the Ch-E solution is computed; we shall work to second order [44, 45, 46], which in the Anderson-Witting case returns the theory already analyzed in [47].
@@
Once the parameterization has been chosen, the next step is to find equations of motion for the $C_{n}$ parameters. We require that these equations both conserve energy-momentum and enforce the Second Law. Note that even if the kinetic theory allows for a $H$-theorem, a positive entropy production in the hydrodynamic theory does not follow automatically, because the parameterized 1-pdf is not a solution of the kinetic equation. As shown in [40], a suitable set of equations of motion is derived by taking the moments of the kinetic equation against the $X^{\alpha}$ functions themselves, see below, eq. (88).
Having found the hydrodynamic equations, we shall consider linearized deviations from equilibrium and derive the propagation speeds [38, 39] for scalar, vector and tensor perturbations. We define the propagation speed as the velocity of a front across which the $C_{\alpha}$ parameters are continuous but their first derivatives are not.
@@
The propagation speeds of a theory are of course fundamental to determine causality. The theories we are considering are thermodinamically stable by construction, and our results will confirm the expectation that they are causal as well [48, 49, 50, 51, 52, 53].
@@
Propagation speeds are also relevant for the discussion of shocks [38, 54, 55]. The propagation speed in kinetic theory is the velocity of the fastest particle for which the 1-pdf is not zero, and so it can be arbitrarily close to the speed of light $c$ for a suitable 1-pdf [56]. Hydrodynamics, on the other hand, usually has a fastest propagation speed which is less than $c$ by a finite amount [39]. For this reason, strong enough shock in hydrodynamics are discontinuous. This discontinuity is not observed in kinetic theory, and may be regarded as an artifact of the hydrodynamic approximation. This makes the issue of which hydrodynamic setup yields the fastest speed most relevant, as this is also the framework which provides the best description of shocks.
@@
To make the discussion more concrete we shall consider a particular family of generalized RTA’s where the relaxation time takes the form $\tau=\tau_{0}(-\beta_{\mu}p^{\mu})^{-a}$, with $-\infty<a<2$, where $\beta_{\mu}=u_{\mu}/T$ is the temperature vector in the Landau frame of the fluid. This family covers both the case where hard modes thermalize faster than soft modes and the converse. It also contains Marle’s and Anderson-Witting RTA’s as the $a=0$ and $a=1$ particular cases, respectively. The upper limit in $a$ is necessary to avoid infrared divergences in the equations of motion.
We shall show that the hierarchy of scalar perturbations propagating faster than vector perturbations, and these being faster than tensor perturbations is upheld [39], and that the Anderson-Witting choice $a=1$, where we recover the results of [47], yields the fastest speeds.
@@
To summarize, the main results of this paper are a) the construction of a generalized RTA designed to match the spectrum of any linearized kinetic equation, b) the derivation thereby of a hydrodynamic theory which is causal and stable and enforces both energy-momentum conservation and positive entropy production, c) the computation of the propagation speeds for scalar, vector and tensor perturbations away from equilibrium for a family of generalized RTA’s containing Marle’s and Anderson-Witting’s as particular cases, and d) the verification that Anderson-Witting RTA yields the fastest speed within this family.
@@
II Kinetic theory
The central object of a kinetic description is the one-particle distribution function (1-pdf) which gives the probability of finding a particle within a given phase space cell, at a particular event and with a particular momentum, constrained to be on mass shell and to have positive energy [57, 58, 59, 60]. For simplicity we shall consider only gases whose equilibrium distribution is of the Maxwell-Jüttner kind eq. (3). The 1-pdf is advected by the particles and changes because of collisions among particles. Therefore the kinetic equation has a transport part and a collision integral which gives the change in the 1-pdf due to collisions per unit particle proper time. There is fairly universal agreement about the transport part, while different kinetic approaches posit different collisions operators [16].
Generally speaking, nonlinear collision operators such as Boltzmann’s lead to very complex theories, and so in practice many times linearized collision integrals are chosen instead. A collision integral must be consistent with energy momentum conservation (for simplicity, we shall deal below with a conformal fluid, and thus we do not impose particle number conservation) and allow for an $H$ theorem, entropy production being zero only for Maxwell-Jüttner 1-pdfs. Concretely, we shall call $f\left(x,p\right)$ a generic 1-pdf, where $p^{2}=0$ and $p^{0}\geq 0$, and
@@ II Kinetic theory
an equilibrium Maxwell-Jüttner distribution for massless particles, where $\beta_{\mu}=u_{\mu}/T$, with $T$ being the fluid temperature and $u_{\mu}$ the velocity, in a frame to be chosen below. We adopt $+$-$+$- + + + signature so $u^{2}=-1$. The kinetic equation reads
$p^{\mu}\frac{\partial f}{\partial x^{\mu}}=\mathcal{I}_{coll}$ (4)
where $\mathcal{I}_{coll}\left[f_{0}\right]=0$. The energy momentum tensor and entropy flux are
@@
$\int Dp\;\ln f\;\mathcal{I}_{coll}\leq 0$ (9)
so that the entropy production $\sigma=S^{\mu}_{,\mu}\geq 0$ for any solution of the kinetic equation eq. (4).
Close to equilibrium we may seek a solution by linearizing around the 1-pdf eq. (3)
$f=f_{0}\left(1+\chi\right)$ (10)
Thus we get
@@
where

### v2

# Propagation Speeds of Relativistic Conformal Particles from a Generalized Relaxation Time Approximation

@@
Propagation Speeds of Relativistic Conformal Particles from a Generalized Relaxation Time Approximation
Alejandra Kandus
LATO-DCET Universidade Estadual de Santa Cruz,Ilhéus-BA, Brazil
Esteban Calzetta
Universidad de Buenos Aires, Facultad de Ciencias Exactas y Naturales,
Departamento de Física, Argentina,
y CONICET - Universidad de Buenos Aires,
Instituto de Física de Buenos Aires (IFIBA), Argentina
kandus@uesc.brcalzetta@df.uba.ar
Resumen
The propagation speeds of excitations are a crucial input in the modeling of interacting systems of particles. In this paper we assume the microscopic physics is described by a kinetic theory for massless particles, which is approximated by a generalized relaxation time approximation (RTA) where the relaxation time depends on the energy of the particles involved. We seek a solution of the kinetic equation by assuming a parameterized one-particle distribution function (1-pdf) which generalizes the Chapman-Enskog (Ch-En) solution to the RTA. If developed to all orders, this would yield an asymptotic solution to the kinetic equation; we restrict ourselves to an approximate solution by truncating the Ch-En series to second order. Our generalized Ch-En solution contains undetermined space-time dependent parameters, and we derive a set of dynamical equations for them applying the moments method. We check that these dynamical equations lead to energy-momentum conservation and positive entropy production. Finally, we compute the propagation speeds for fluctuations way from equilibrium from the linearized form of the dynamical equations. Considering relaxation times of the form $\tau=\tau_{0}(-\beta_{\mu}p^{\mu})^{-a}$, with $-\infty<a<2$, where $\beta_{\mu}=u_{\mu}/T$ is the temperature vector in the Landau frame, we show that the Anderson-Witting prescription $a=1$ yields the fastest speed in all scalar, vector and tensor sectors. This fact ought to be taken into consideration when choosing the best macroscopic description for a given physical system.
1 Introduction
The success of hydrodynamics in the description of the early stages of relativistic heavy ion collisions [1, 2] and the promise of relevant cosmological applications [3, 4, 5] have turned the study of strongly interacting systems of relativistic particles into an active area of research [7, 8, 6]. Although kinetic theory provides the microscopic description for these systems [9, 10, 11, 12], the full Boltzmann equations are generally regarded as too complex, and simpler schemes consistent with the most important physical features are sought. Among these simpler schemes, the Relaxation Time Approximation (RTA), which assumes a collision term parameterized by a relaxation time, stands out. The RTA includes the non linearities in the theory through the equilibrium one particle distribution function (1-pdf) towards which the system relaxes, and in this sense the RTA ansatz contains the footprint of the non-linear and complete underlying kinetic theory that describes the microphysics.
The first implementations of the Bhatnagar, Gross and Krook BGK RTA [13] to relativistic fluids are those of Marle [14, 15] and Anderson and Witting (AW) [16, 17]. Both expressions are phenomenological ansätze, proportional to $(f-f_{0})$ and in each case $f_{0}$ corresponds to different ’frames’ (see sub-section 3.3 below). In the case of Marle, the proportionality factor is $m/\tau$, with $m$ the mass of the gas particles, while in that of AW it is $u_{\mu}p^{\mu}/\tau$. In both cases the parameter $\tau$ is identified with a ’relaxation time’. In principle the choice of the proportionality factor must be guided by phenomenological considerations or else by a systematic derivation form the Boltzmann equation, which is feasible in some cases. See Refs. [18, 19]. However successful, both Marle and AW RTA’s severely distort the structure of the linearized Boltzmann equation, and their validity is doubtful for soft collision terms, where the continuous spectrum of the linearized collision operator reaches up to the zero eigenvalue associated with the hydrodynamic modes. [20, 21, 22].
@@ 1 Introduction
This has led to more general implementations of the RTA [23, 24] where different modes were allowed to relax at different rates. However, simply allowing the relaxation times to be a function of energy other than constant or linear is not satisfactory as it may violate energy-momentum conservation [25, 26]. Rather the RTA must be implemented preserving the Hilbert space structure of the space of linearized one-particle distribution function (1-pdf’s). A kinetic equation allowing for a momentum dependent relaxation time consistent with energy-momentum conservation was introduced in [27, 28]. In this paper we shall elaborate on this proposal. In particular, we shall show how to produce a RTA matching any prescribed spectrum for the linearized collision operator, either soft or hard. Not been able to do this is one of the main drawbacks of the usual formulations of the RTA. See also [29, 30, 31, 32, 33, 34, 35, 36, 37, 38]
Although a general solution of the RTA kinetic equation may be attempted, this leads to an integral equation for the equilibrium 1-pdf, which must be related to the actual 1-pdf through some prescription, which will be different for different choices of the relaxation time. For example, under the Marle prescription the equilibrium 1-pdf leads to the same particle current than the actual one, while under the Anderson-Witting prescription, the equilibrium and actual 1-pdfs are matched though the energy current. This obscures the physical features of the system, which close enough to equilibrium is dominated by the hydrodynamic modes and a few long-lived non-hydrodynamic modes. To capture this behavior, it is best to assume a parameterized form for the 1-pdf, supplemented by dynamical equations for the parameters. In this paper, we shall obtain these dynamical equations by taking moments of the kinetic equation.
Of course this poses the challenge of finding a suitable parameterization for the 1-pdf. The parameterization should be general enough to allow for an accurate description of physically meaningful processes but not so general as to make the ensuing theory unwieldly. Many proposals have been advanced in the literature [39, 40, 41, 42, 43, 44] .
In this paper we shall adopt the point of view that the parameterization must be such as to include the Chapman-Enskog (Ch-En) solution to the kinetic equation as a particular case [45]. In other words, we shall use the second order Ch-En as a template, generalizing it so to obtain a family of parameterized 1-pdf, still containing the actual Ch-En solution as a particular case.
One reason for working this way is the fact that in the Ch-En expansion, as in DNMR [46] and IReD [47], each order is determined by the power of a certain small parameter. Thus it is easy to identify the intensity of the deviation from the local thermal balance.
The procedure we propose yields parameterizations with an increasing complexity depending on to which order the Ch-En solution is computed. In this manuscript we shall work to second order [48, 49, 50], which in the AW case returns the theory already analyzed in [51].
@@
Once the parameterization has been chosen, the next step is to find equations of motion for the parameters. We require that these equations both conserve energy-momentum and enforce the Second Law. Note that even if the kinetic theory allows for a $H$-theorem, a positive entropy production in the parameterized theory does not follow automatically, because the parameterized 1-pdf is not a solution of the kinetic equation. As shown in [52], a suitable set of equations of motion is derived by taking the moments of the kinetic equation against the same irreducible phase-space functions which appear in the parameterization, see below, eq. (17). Observe that in principle this method does not require linearization on deviations from equilibrium (see also [53]), though in general the resulting theory is too complex unless severely restricted by symmetry considerations.
The ultimate goal of this paper is to compute the propagation speeds for collective modes of a conformal real relativistic system of interacting particles [54, 55, 56, 57, 58, 59]. We define the propagation speed as the velocity of a front across which the parameters are continuous but their first derivatives are not. To derive the hydrodynamic equations we apply the moments method to a kinetic equation under the RTA. We work in the Landau frame of the fluid, and generalize the usual RTA by allowing an arbitrary dependence of relaxation time on the particle energy [27, 28]. Different choices for this dependence mean different underlying kinetic theories or equivalently, microphysics.
One crucial step in this procedure is the choice of which moments of the 1-pdf shall be considered. As we said above, we adopt the criterium of choosing the functions of momentum which make up a second order Chapman-Enskog solution. In other words, we work within a restricted class of 1-pdfs which generalizes the second order Chapman-Enskog solution, retaining it as a particular case. This choice leads us to parametrize the 1-pdf in terms of functions of momentum which themselves depend on the way the relaxation time relates to energy, see eqs. (65) below. Because of this, different choices of the functional dependence of the relaxation time lead to different dynamical equations for the system, not only in the terms which derive from the moments of the collision integral, where the dependence on the relaxation time is explicit, but also in the derivative terms. We thus obtain a family of dynamical theories with different phenomenology according to the choice of the relaxation time as a function of energy. We could understand this dependence as the remaining footprint of a non-linear and complete kinetic theory, which was approximated by a mathematically more tractable RTA.
In summary, assuming the relaxation time dependent upon energy [25, 26, 27, 28] we show, on one hand, how to compare the predictions from different choices of this dependence, and, on the other, that those different choices lead to macroscopic models that are clearly distinguishable (in this case by producing different propagation speeds).
@@
The propagation speeds of a theory are of course fundamental to determine causality. The theory we are considering here is thermodinamically stable by construction, and our results will confirm the expectation that it is causal as well [60, 61, 62, 63, 64, 65, 66].
@@
Propagation speeds are also relevant for the discussion of shocks [54, 67, 68]. The propagation speed in kinetic theory is the velocity of the fastest particle for which the 1-pdf is not zero, and so it can be arbitrarily close to the speed of light $c$ for a suitable 1-pdf [69]. There are also examples from field theory where the propagation speed is arbitrarily close to that of light [70]. Hydrodynamics, on the other hand, usually has a fastest propagation speed which is less than $c$ by a finite amount [58]. For this reason, strong enough shocks in hydrodynamics are discontinuous. This discontinuity is not observed in kinetic theory, and may be regarded as an artifact of the hydrodynamic approximation. When considering approximations to the full kinetic equations, as in this paper, the issue of which setup yields the fastest speed becomes most relevant, as this is also the framework which provides the best description of shocks.
@@
To make the discussion more concrete we shall consider a particular family of generalized RTA’s where the relaxation time takes the form $\tau=\tau_{0}(-\beta_{\mu}p^{\mu})^{-a}$, with $-\infty<a<2$, where $\beta_{\mu}=u_{\mu}/T$ is the temperature vector in the Landau frame of the fluid. This family covers both the case where hard modes thermalize faster than soft modes and the converse. It also contains Marle’s and AW RTA’s as the $a=0$ and $a=1$ particular cases, respectively. The upper limit in $a$ is necessary to avoid infrared divergences in the equations of motion.
We find that the AW choice $a=1$, where we recover the results of [51], yields the fastest speeds.
@@
To summarize, the main results of this paper are a) the construction of a generalized RTA designed to match the spectrum of any linearized kinetic equation, b) the derivation thereby of a parameterized theory which is causal and stable and enforces both energy-momentum conservation and positive entropy production, c) the computation of the propagation speeds for scalar, vector and tensor perturbations away from equilibrium for a family of generalized RTA’s containing Marle’s and AW’s as particular cases, and d) the verification that AW RTA yields the fastest speed within this family.
This paper is organized as follows: In Section 2 we shortly review the features of Kinetic Theory, and in section 3 we elaborate on the generalized RTA as a substitute for the actual kinetic equation as derived from microphysics. In Section 4 we deduce the 1-pdf Ch-En solution to the Boltzmann equation up to second order in gradients, for a momentum dependent RTA. We then introduce a 1-pdf after the pattern of the Ch-En solution and derive the set of moment equations, which, for the purposes of finding the propagation speeds, may be particularized at the free-streaming regime. In Section 5 we perform a scalar-vector-tensor decomposition and write down the equation system corresponding to each sector. To have a glimpse of their solutions we consider the family of RTA’s given by $\tau=\tau_{0}(-\beta_{\mu}p^{\mu})^{-a}$, with $-\infty<a\leq 2$ to avoid infrared divergences. In Section 6 we summarize the main conclusions. Details of the derivation of the Ch-En solution are given in the appendix A.
@@
2 Relativistic Kinetic theory
The central object of a kinetic description is the one-particle distribution function (1-pdf) which gives the probability of finding a particle within a given phase space cell, at a particular event and with a particular momentum, constrained to be on mass shell and to have positive energy [71, 72, 73, 74]. For simplicity we shall consider only gases whose equilibrium distribution is of the Maxwell-Jüttner kind eq. (2). The 1-pdf is advected by the particles and changes because of collisions among particles. Therefore the kinetic equation has a transport part and a collision integral which gives the change in the 1-pdf due to collisions per unit particle proper time, as in
$p^{\mu}\frac{\partial f}{\partial x^{\mu}}=\mathcal{I}_{coll}$ (1)
There is fairly universal agreement about the transport part, while different kinetic approaches posit different collisions operators [6]. A collision integral must be consistent with energy momentum conservation (for simplicity, we shall deal below with a massless particles, and thus we do not impose particle number conservation) and allow for an $H$ theorem, entropy production being zero only for Maxwell-Jüttner 1-pdfs
@@ 2 Relativistic Kinetic theory
where $\beta_{\mu}=u_{\mu}/T$, with $T$ being the fluid temperature and $u_{\mu}$ the velocity in a frame to be chosen below, with $u^{2}=-1$. Then
$\mathcal{I}_{coll}\left[f_{0}\right]=0$ (3)
The energy momentum tensor (EMT) and entropy flux are
@@
$-\int Dp\;\ln f\;\mathcal{I}_{coll}\geq 0$ (8)
for any function $f$. If $f$ is a solution of the kinetic equation eq. (1) this leads to positive entropy production $S^{\mu}_{,\mu}\geq 0$.
The Landau-Lifshitz prescription provides a way to associate an inverse temperature vector to any 1-pdf, even if it is not of the Maxwell-Jüttner kind. Namely, we identify the four-velocity with the only timelike eigenvector of the EMT
$T^{\mu}_{\nu}u^{\nu}=-\rho u^{\mu}$ (9)
whereby we identify $\rho$ as the energy density, and then we derive a temperature from $\rho$ by imposing the equilibrium dependence
$\rho=\frac{3}{\pi^{2}}T^{4}$ (10)
Having identified $\beta^{\mu}$, we may build the corresponding Maxwell-Jüttner 1-pdf $f_{0}$, eq. (2). Moreover, the fact that $\ln f$ appears explicitly in the H-theorem eq. (8) suggests we decompose $f$ as
$f=f_{0}e^{\chi}$ (11)
Without loss of generality we may also write
@@
By virtue of eq. (7) $\ln f_{0}$ does not contribute to entropy creation, and then we find
$S^{\mu}_{;\mu}=-\int Dp\;f_{0}\;\chi\;{I}_{coll}\left[\chi\right]\geq 0$ (13)
If $f$ is a solution of the kinetic equation.
The ansatz eq. (11) guarantees a positive one-particle distribution function, even in a full nonlinear theory. Moreover, the entropy production eq. (13) already singles out the logarithm of the one-particle distribution function as playing a most important role. Using this feature we will introduce below an ansatz for the collision term (see eq. (34 )) which leads to positive entropy production to all orders in deviations from equilibrium, (see eq. (35)).
2.1 Parameterized Kinetic Theory
Let us look for solutions of eq. (1) of the form
$f\equiv f_{hydro}=e^{\sum_{\alpha=0}^{n}C_{\alpha}X^{\alpha}}$ (14)
In eq. (14), the $C_{\alpha}$ denote tensor fields in space-time, while the $X^{\alpha}$ are tensor fields in phase space. Note that $\alpha$ is not itself a tensorial index, it just numbers the different tensors in the theory. The contractions $C_{\alpha}X^{\alpha}$ are world scalars. In particular, we shall choose $X^{0}=p^{\mu}$ and $C_{0}=\beta_{\mu}$, or equivalently, in the terms of eq. (11)
$\chi=\sum_{\alpha=1}^{n}C_{\alpha}X^{\alpha}$ (15)
We shall assume $C_{\alpha}\left(x\right)$ are totally symmetric, traceless and transverse tensors, for $\alpha\geq 1$. Of course, if we would allow the $X^{\alpha}$ functions to form a complete set in phase space we could seek an exact solution of the kinetic equations under this form. We shall however truncate the sum in eq. (15) at a finite value of $n$, to be specified below, so we shall get only an approximate solution.
The EMT and entropy flux are obtained by substituting $f_{hydro}$ into equations (4) and (5). We obtain the four conservation laws (7) but these equations are not enough to determine the evolution of the whole set of $C_{\alpha}$ functions, for which we must provide supplementary equations.
Our guiding principle is to obtain positive entropy production. Now
$S^{\mu}_{hydro,\mu}=-\int\;Dp\;\left[\sum_{\alpha=1}^{n}C_{\alpha}X^{\alpha}\right]p^{\mu}f_{hydro,\mu}$ (16)
The problem is that we cannot call upon the H-theorem to enforce positive entropy production, because $f_{hydro}$ is not a solution of the kinetic equation. We demand instead the moment equations
$\int Dp\;X^{\alpha}\left\{p^{\mu}f_{hydro,\mu}-f_{0}I_{coll}\left[\chi\right]\right\}=0$ (17)
which for $\alpha=0$ is just EMT conservation. We now are allowed to substitute
$S^{\mu}_{hydro,\mu}=-\int\;Dp\;f_{0}\chi I_{coll}\left[\chi\right]\geq 0$ (18)
The moment equations (17) are thus the equations of motion of the parameterized theory.
This setup enforces positive entropy production but does not tell us how to choose the $X^{\alpha}$ functions, beyond $\alpha=0$. We shall return to this vexing question in section (4), after we have introduced the relaxation time approximation.
3 Relaxation time approximation
Physically, the role of the collision term in the kinetic equation (1) is to force $\chi$ to relax to zero, or at least to a multiple of $p^{\mu}$, the only possibilities leading to vanishing entropy production. A realistic kinetic equation such as Boltzmann’s typically leads to a very complex collision term. However, it may be expected that the essentials of the relaxation of $\chi$ may be captured by a much simpler collision term, linear in $\chi$